OpenStudy (anonymous):

Can I get help on how to find the vertex for an equation? ex: y=2x^-4x-5

OpenStudy (tkhunny):

I presume you mean \(y = 2x^{2} - 4x - 5\) Two ways. One you either know or you don't. The x-coordinate of the vertex is often stated as -b/2a. You have a = 2 and b = -4, giving x = 1 and the corresponding y = -7. Otherwise, how's your "Completing the Square"? \(y = 2(x^{2} - 2\;+\;??) - 5 - 2(??)\) -- Fill in the blanks.

OpenStudy (anonymous):

-4?

OpenStudy (tkhunny):

-2/2 = -1 (-1)^2 = 1 \(y=2(x^2−2x+1)−5−2(1)\) <== Whoops!! I dropped the 'x' up above. Sorry about that. \(y=2(x-1)^{2}−7\) and you can just read off the Vertex. (1,-7)

OpenStudy (anonymous):

thanks..